14. A criminal is escaping across a rooftop and runs off the roof horizontally at a speed of 5.3 m/s, hoping to land on the roof of an adjacent building. Air resistance is negligible. The horizontal distance between the two buildings is D, and the roof of the adjacent building is 2.0 m below the jumping-off point. Find the maximum value for D.
Answer #1

initial vertical velocity=0

vertical acceleration=g=9.8

vertical displacement=2 m

let time taken=t

s=u*t+(1/2)*g*t2

2=0+(1/2)*9.8*t2

t=0.64 s

initial horizontal velocity=5.3

horizontal acceleration=0

t=0.64

D=u*t+(1/2)*a*t2

D=5.3*0.64+0

D=3.39m (ans)